A cone of radius π‘Ÿ centimeters and height β„Ž centimeters is lowered point first at


a rate of 1 cm/s into a tall cylinder of radius 𝑅 centimeters that is partially filled with


water. How fast is the water level rising at the instant the cone is completely


submerged

1
Expert's answer
2022-04-05T16:37:33-0400

Potential energy is converted into kinetic energy of translational and rotational motion.

"m*g*h=m*v\u00b2\/2 + J*\u03c9\u00b2\/2"

moment of inertia of the cylinder

"J=m*r\u00b2\/2"

Angular velocity at the end of the inclined plane

"ω=v/r\, \\

m*g*h=m*v²/2 + (m*r²/2)*(v²/(2*r²))\\

g*h=v²/2 + v²/4=(3/4)*v²\\

v=2*\sqrt{g*h/3}"

Answer: "v=2*\sqrt{g*h/3}"